Goldberg Corporation has three production departments A, B, and C. Goldberg Corporation also has two service departments, Administration and Personnel. Administration costs are allocated based on value of assets employed, and Personnel costs are allocated based on number of employees. Assume that Administration provides more service to the other departments than does the Personnel Department.
Dept. Direct Costs Employees Asset Value
Admin. $900,000 25 $450,000
Personnel 350,000 10 600,000
A 700,000 15 300,000
B 200,000 5 150,000
C 250,000 10 800,000
Refer to Goldberg Corporation. Using the step method, what amount of Administration costs is allocated to A (round to the nearest dollar)? Show work.
Answer #1

Total asset value for allocation = 600000+300000+150000+800000 = 1850000

Administration cost allocated to A = 900000/1850000*300000 = $145946